OpenStudy (anonymous):

What is the vertex for the graph of y = 2x2 + 12x + 10?

OpenStudy (anonymous):

2x^2

OpenStudy (jdoe0001):

you can find the vertex for a vertical parabola of that type by using => \(\bf \left(-\cfrac{b}{2a}, c-\cfrac{b^2}{4a}\right)\)

OpenStudy (anonymous):

I use the ^ to indicate the exponent. So the equation would be written: y = 2x^2 + 12x + 10 = 2 (x + 1) (x + 5) so this parabola intersects the x axis at x=-1 and x=-5. The midpoint between them is the x-coordinate of the vertex: [(-1) + (-5)] / 2 = -3 and the y-coordinate is given by the function at that point: y = 2 * (-3)^2 + 12 (-3) + 10 = 18 - 36 + 10 = -8 The vertex is (-3, -8).

OpenStudy (cwrw238):

another way to find the vertex is to convert the function to vertex form 2x^2 + 12x + 10 = 2 (x^2 + 6x + 5) = 2((x + 3)^2 - 9 + 5) = 2 (x + 3)^2 - 4) = 2 (x + 3)^2 - 8 the vertex is at the point where (x +3) = 0 i.e. x = -3 and y = -8 vertex is (-3, -8)